OpenStudy (anonymous):

write an expression for the perimeter in simplest form. it is a rectangle (3x-2)(3x-2)+(5x+7)(5x+7) my answer is 16x+10 is that right

OpenStudy (radar):

I think you're right, but you have them expressed as a combination of multiplication and addition. I assume ones side is 3x-2, and the other dimension is 5x+7. The perimeter would be expressed as P=2(3x-2) + 2(5x+7) P=6x-4 + 10x+14)=16x +10

OpenStudy (radar):

If I followed the way you expressed it, it would be\[(9x ^{2}-12x+4)+(25x ^{2}+70x +49)\] A totally different looking animal.

OpenStudy (radar):

Do you see the difference?